OpenStudy (anonymous):

Three boys and three girls are to sit together in a row. Find the probability that : (a) the boys and girls alternate. (b) the boys and girls sit together. (c) two specific girls sit next to one another

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok? lol

OpenStudy (anonymous):

can anyone help me out?

OpenStudy (nilankshi):

good

OpenStudy (anonymous):

good what?

OpenStudy (nilankshi):

every thing is good

OpenStudy (anonymous):

if someone explains this question everything will be good :)

OpenStudy (nilankshi):

hahaha

OpenStudy (nilankshi):

bye

OpenStudy (anonymous):

bye nilankshi

OpenStudy (nilankshi):

bye

OpenStudy (anonymous):

joe helllp me lol

OpenStudy (anonymous):

is it 2x (3! 3!/6!)

OpenStudy (anonymous):

are we able to tell between people? like, are these people b1, b2, b3, g1, g2, g3? or b, b, b, g, g, g, where we cant tell between them.

OpenStudy (anonymous):

you can't tell

OpenStudy (anonymous):

i dont like these problems because its never too clear.

OpenStudy (anonymous):

lol i hate probability in general

OpenStudy (anonymous):

my answer is right but i don't know if my method is right

OpenStudy (anonymous):

Well, there are 2 ways we can have boys and girls alternate, b g b g b g or g b g b g b so only 2. There are\[\frac{6!}{3!3!}=\frac{6\cdot 5\cdot 4}{1\cdot 2 \cdot 3}=20\]total combinations, so the probability is: \[\frac{2}{20}=\frac{1}{10}\] Which i believe is what you got for the first answer.

OpenStudy (anonymous):

There are only 2 ways that boys can sit together and girls can sit together, namely: b b b g g g or g g g b b b so the probability will be the same again, 1/10

OpenStudy (anonymous):

what is the formula again?

OpenStudy (anonymous):

For the total number of combinations?

OpenStudy (anonymous):

because my working was the other way around to yours

OpenStudy (anonymous):

I dont really know formulas =/ I just know that since there are 6 people, the number of ways their order can be changed is 6! This is because when i pick someone to be the first person, I have 6 people to choose from. When I pick the second person there will be 5 to choose from. Then 4, then 3, and so on, so the total number of combinations is: 6*5*4*3*2*1=6! But because i cant tell the difference between the boys if i switch them around, or the girls, I over counted. I have to divide by the number of ways I can switch the boys only, and the number of ways i can switch the girls only. So i take that 6!, and divide by 3! for the boys, and again by 3! for the girls.

OpenStudy (anonymous):

isee thanks for the explanation :)

OpenStudy (anonymous):

so c ) will be 6!/4!2!

OpenStudy (anonymous):

yes . it's 6P2 = 6!/4! 2!